OpenStudy (anonymous):

g(x)=x^2-10x-3 a) g(x+5) b) g(-x)

OpenStudy (zzr0ck3r):

g(x) = x^2-10x-3 g(3) means put a 3 where you see an x so g(x+5) means put an x+5 where you see an x

OpenStudy (anonymous):

i have for a) x^2+72

OpenStudy (anonymous):

is that true?

OpenStudy (zzr0ck3r):

x^2-10x-3 so (x+5)^2-10(x+5)-3 x^2+10x+25-10x-50-3 x^2+25-50-3 x^2-28

OpenStudy (anonymous):

If you have \[g(x) = x ^{2} - 10x -3\] and you want to find g(x+5) you simply replace every 'x' in your original equation with an 'x+5' So it now becomes \[g(x+5) = (x+5)^{2} -10(x+5)-3\] and simplify the same for g(-x) \[g(-x)= (-x ^{2}) -10*(-x)-3\] [don't forget the square of a negative is positive!] Hope that helped!

OpenStudy (anonymous):

when i get this far on a) x^2+10x+25-10x-50-3 next step is to order terms before any math is completed?